# Test-Time Training Provably Improves Transformers as In-context Learners

### v1

@@ Appendix B Proofs for Section 4
Recall the standard result, which follows from Wick’s Theorem (Petersen & Pedersen, 2012)[Section 8.2.4] for a zero-mean Gaussian $\bm{x}\sim\mathcal{N}(\bm{0},\bm{\Sigma})$ and any symmetric matrix ${\bm{A}}$:
@@
Hence, considering that we have $n$ diagonal terms and $n(n-1)$ off-diagonal terms, combining (i) + (ii) gives:
@@
${\cal{L}}(\bm{W})$ $={\bm{\beta}_{\text{TT}}}^{\top}\left[\bm{\Sigma}_{\bm{x}}\;-\;n\bm{\Sigma}_{\bm{x}}\,\bm{W}\,\bm{\Sigma}_{\bm{x}}\;-\;n\bm{\Sigma}_{\bm{x}}\bm{W}^{\top}\,\,\bm{\Sigma}_{\bm{x}}\;+\;n(n+1)\bm{\Sigma}_{\bm{x}}\bm{W}^{\top}\bm{\Sigma}_{\bm{x}}\bm{W}\bm{\Sigma}_{\bm{x}}+n\texttt{tr}(\bm{W}^{\top}\bm{\Sigma}_{\bm{x}}\bm{W}\bm{\Sigma}_{\bm{x}})\bm{\Sigma}_{\bm{x}}\right]{\bm{\beta}_{\text{TT}}}$
@@
We consider the noiseless ($\sigma^{2}=0$) and isotropic case where $\bm{\Sigma}_{\bm{x}}=\bm{\Sigma}_{\bm{\beta}}={\bm{I}}$. Throughout the proof, we omit lower order terms like $+1$ in $(d+1)$ or $2n$ in $(n^{2}+2n)$ as they’re negligible compared to the higher order of the same variable (see Remark B.4 for error discussion). This convention allows us to simplify the subsequent expressions.
@@
A key technical challenge while using the above update is that the expectation of $\Delta{\cal{L}}$ in (10) involves $8th$ moments of ${\bm{X}_{\text{context}}}$. To overcome this problem, we will use the Gaussian approximation following Lemma B.1 by approximating ${\bm{X}_{\text{context}}}^{\top}{\bm{X}_{\text{context}}}\,{\bm{\beta}_{\text{TT}}}\approx n\bm{\Sigma}_{\bm{x}}\,{\bm{\beta}_{\text{TT}}}+\sqrt{n}\,\bm{\Sigma}_{\bm{x}}^{1/2}{\bm{g}}$ where ${\bm{g}}\sim\mathcal{N}(\bm{0},\|{\bm{\beta}_{\text{TT}}}\|^{2}\,{\bm{I}})$. This non-isotropic form can be obtained by factoring out $\bm{\Sigma}_{\bm{x}}^{1/2}$ and reducing the problem to the isotropic case from Lemma B.1. Also define $\Delta\bm{W}:={\bm{W}_{\text{TT}}}-\bm{W}^{*}$. Then, using the approximation and the fact that $\bm{\Sigma}_{\bm{x}}={\bm{I}}$, we can write:
@@
$=4\eta^{2}\frac{k(k+d+1)n}{(n+d+1)^{2}}\operatorname{\mathbb{E}}_{{\bm{g}}}\left[\left(\sqrt{n}(d+1)\,{\bm{\beta}_{\text{TT}}}\,{\bm{\beta}_{\text{TT}}}^{\top}+(d+1){\bm{g}}\,{\bm{\beta}_{\text{TT}}}^{\top}-n\,{\bm{\beta}_{\text{TT}}}{\bm{g}}^{\top}-\sqrt{n}{\bm{g}}{\bm{g}}^{\top}\right)\left(\sqrt{n}(d+1)\,{\bm{\beta}_{\text{TT}}}\,{\bm{\beta}_{\text{TT}}}^{\top}\right.\right.$
@@
$=-4\eta kn(n+1)\frac{nd}{(n+d+1)^{2}}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4}$
$\quad-4\eta^{2}\frac{k(k+d+1)n^{2}(n+1)}{(n+d+1)^{2}}\left[n\left(d^{2}-2d-3+nd\right)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{6}+\left((d+1)^{2}+n(d+2)\right)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{6}\right],$ (14)
@@
${\cal{L}}(\bm{W}^{*})-{\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$
$\quad=4\eta k\frac{n^{2}d}{n+d+1}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4}-4\eta k\frac{n^{2}}{(n+d+1)^{2}}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4}-4\eta k\frac{n^{2}(n+1)d}{(n+d+1)^{2}}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4}$
$\qquad-4\eta^{2}\frac{k(k+d+1)n^{2}}{(n+d+1)^{2}}\left[n(n+2)(d^{2}-2d-3+nd)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{6}+(n+d+1)\left((d+1)^{2}+n(d+2)\right)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{6}\right].$
@@
Taking the derivative of this expression and setting to 00 yields that the optimal $\eta^{*}$ solution is approximately:
@@
Remark B.4.
We note that in the proportional $n,d$ regime, the Gaussian approximation in Lemma B.1 introduces an error of $\mathcal{O}(n^{-1})$. Additionally, when we drop lower-order terms (such as $2n$ in $n^{2}+2n$), the corresponding error is of one degree lower than the original expression. Consequently, after carrying these approximations through the entire derivation, the overall error remains $\mathcal{O}(n^{-1})$, since the final loss expression is zeroth-order in $n,d$. Hence, our result is robust to these approximations as $n,d\rightarrow\infty$ while maintaining the ratio $\alpha=n/d$. The same reasoning applies to the non-isotropic covariance case as well.
See 4.3
@@
Using Theorem 4.2, the new loss ${\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$ is approximately:
@@
[…]
@@
$\operatorname{\mathbb{E}}_{{\bm{X}_{\text{train}}},{\bm{\xi}_{\text{train}}},{\bm{X}_{\text{context}}},{\bm{\xi}_{\text{context}}}}\left[{\bm{\beta}_{\text{TT}}}^{\top}\,n\,\bm{\Sigma}_{\bm{x}}\,\left({\bm{W}_{\text{TT}}}-\bm{W}^{*}\right)\,\bm{\Sigma}_{\bm{x}}\,{\bm{\beta}_{\text{TT}}}\right]$ (20)
$\qquad=2\eta n\,{\bm{\beta}_{\text{TT}}}^{\top}\operatorname{\mathbb{E}}_{{\bm{X}_{\text{train}}},{\bm{X}_{\text{context}}}}\left[{\bm{X}_{\text{train}}}^{\top}{\bm{X}_{\text{train}}}\,{\bm{\beta}_{\text{TT}}}\,{\bm{\beta}_{\text{TT}}}^{\top}{\bm{X}_{\text{context}}}^{\top}{\bm{X}_{\text{context}}}\right]{\bm{\beta}_{\text{TT}}}$
$\qquad=2\eta kn^{2}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4},$ (21)
@@
$=4kn^{2}\left[\eta\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4}-\eta^{2}\left(\sigma^{2}d+(k+d+1)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}\right)\left[\left(\sigma^{2}+\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}\right)\left(\sigma^{2}d+(n+d+1)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}\right)\right.\right.$
@@
Setting the derivative to 00 and solving for $\eta$ gives:
@@
In the proportional $n,d$ regime and when $k/d\to\infty$, this gives an error of order $\mathcal{O}(n^{-1})$-matching that of the optimal ${\bm{W}^{\text{opt}}}$ from Lemma C.1:
@@
Considering that the initial loss of the weight matrix $\bm{W}=\bm{0}_{d\times d}$ is $\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}$ and the improvement given by Theorem 4.4, the new loss after the test-time-training update is approximately:
@@
On the other hand, recall that by Proposition 4.1, the loss of $\bm{W}^{*}$ before the update is $\frac{d+1}{n+d+1}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}$. Combining this with the improvement given by Theorem 4.2, the new loss after the test-time-training update is approximately:
@@ Appendix C Proofs for Section 5
[…]
@@
Hence, all other eigenvalues of $\bm{W}$ are 00. Using the eigendecomposition of $\bm{W}$, this uniquely specifies $\bm{W}$ as $c\cdot{\bm{\beta}_{\text{TT}}}\,{\bm{\beta}_{\text{TT}}}^{\top}$. Solving for $c$ yields the unique solution to $\nabla_{\bm{W}}{\cal{L}}(\bm{W})=0$:
@@
Consider the optimal pre-trained $\bm{W}^{*}$ matrix, which minimizes the population loss over all possible tasks sampled from $\bm{\Sigma}_{\bm{\beta}}$ in (1). Then, all eigenvalues of $\bar{\bm{W}}^{*}=\bm{\Sigma}_{\bm{x}}^{1/2}\bm{W}^{*}\bm{\Sigma}_{\bm{x}}^{1/2}$ lie in the interval $[0,\frac{1}{n+1}]$.
@@
preserves the population risks in (1) and (3). More precisely, defining ¯$\bar{{\cal{L}}}(\bar{\bm{W}})$ to be the population loss (3) evaluated under the transformed setting $(\bar{\bm{x}},\bar{{\bm{X}}},\bar{{\bm{\beta}}},\bar{\bm{W}})$, we have
@@
Let $n/d=\Theta(1)$ and $\sigma^{2}=0$. Suppose the covariance matrices are jointly diagonalizable by an orthogonal matrix ${\bm{Q}}\in\mathbb{R}^{d\times d}$, so that $\bm{\Sigma}_{\bm{x}}={\bm{Q}}{\bm{\Lambda}}_{\bm{x}}{\bm{Q}}^{\top}$ and $\bm{\Sigma}_{\bm{\beta}}={\bm{Q}}{\bm{\Lambda}}_{\beta}{\bm{Q}}^{\top}$. Let $\bm{W}$ be any matrix that’s also jointly diagonalizable by ${\bm{Q}}$ such that if $\bm{\Sigma}_{\bm{x}}^{1/2}\,\bm{W}\,\bm{\Sigma}_{\bm{x}}^{1/2}={\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}$, then all eigenvalues of ${\bm{\Lambda}}$ lie in the interval $[0,\frac{1}{n+1}]$. Define ${\tilde{\bm{\beta}}_{\text{TT}}}:={\bm{Q}}^{\top}\bm{\Sigma}_{\bm{x}}^{1/2}{\bm{\beta}_{\text{TT}}}$, $A:={\tilde{\bm{\beta}}_{\text{TT}}}^{\top}\left({\bm{I}}-n{\bm{\Lambda}}\right)^{2}{\tilde{\bm{\beta}}_{\text{TT}}}$, and $B:=n\|{\tilde{\bm{\beta}}_{\text{TT}}}\|_{2}^{2}\texttt{tr}({\bm{\Lambda}}^{2})$. Then, the optimal step size that minimizes the population loss given in (3) after the test-time training update is
@@
As in previous theorems, we will use the Gaussian approximation following Lemma B.1 by approximating ${\bm{X}_{\text{context}}}^{\top}{\bm{X}_{\text{context}}}\,{\bm{\beta}_{\text{TT}}}\approx n\bm{\Sigma}_{\bm{x}}\,{\bm{\beta}_{\text{TT}}}+\sqrt{n}\,\bm{\Sigma}_{\bm{x}}^{1/2}{\bm{g}}$ where ${\bm{g}}\sim\mathcal{N}(\bm{0},\|{\bm{\beta}_{\text{TT}}}^{2}\|\,{\bm{I}})$. This non-isotropic form can be obtained by factoring out $\bm{\Sigma}_{\bm{x}}^{1/2}$ and reducing the problem to the isotropic case from Lemma B.1. Before going forward, let’s apply the covariance shift discussed in Lemma C.3, which transforms:
@@
$=2\eta{\bar{\bm{X}}_{\text{train}}}^{\top}{\bar{\bm{X}}_{\text{train}}}\left(n\,{\bar{\bm{\beta}}_{\text{TT}}}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}+\sqrt{n}\,{\bar{\bm{\beta}}_{\text{TT}}}{\bm{g}}^{\top}-n^{2}{\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}{\bar{\bm{\beta}}_{\text{TT}}}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}-n\sqrt{n}\,{\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}{\bar{\bm{\beta}}_{\text{TT}}}\,{\bm{g}}^{\top}\right.$
$\qquad\left.-n\sqrt{n}\,{\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}{\bm{g}}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}-n{\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}{\bm{g}}{\bm{g}}^{\top}\right)$
@@
Now, recall from Equation 29 that the second-order difference is in the form:
@@
$\approx 4\eta kn^{2}\|{\bar{\bm{\beta}}_{\text{TT}}}\|_{2}^{2}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}\,{\bm{Q}}\left({\bm{I}}-n{\bm{\Lambda}}\right)^{2}{\bm{Q}}^{\top}{\bar{\bm{\beta}}_{\text{TT}}}$
$\quad-4\eta^{2}k(k+d+1)n^{2}\left[n^{3}\,\|{\bar{\bm{\beta}}_{\text{TT}}}\|_{2}^{6}\,\texttt{tr}({\bm{\Lambda}}^{2})+(n^{2}+d)\,\|{\bar{\bm{\beta}}_{\text{TT}}}\|_{2}^{4}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}\,{\bm{Q}}\left({\bm{I}}-n{\bm{\Lambda}}\right)^{2}{\bm{Q}}^{\top}{\bar{\bm{\beta}}_{\text{TT}}}\right].$ (40)
This is a quadratic expression in $\eta$, and we can solve for the optimal $\eta^{*}$ by setting the derivative to 00. This way, the optimal step size is approximately:
@@
While Li et al. (2024) derives the solution form in the full-rank case, the same closed-form expression can be adapted when $\bm{\Sigma}_{\bm{\beta}}$ is low-rank. Indeed, if $\bm{\Sigma}_{\bm{\beta}}$ has zeros in its last $d-r$ diagonal entries, those coordinates do not appear in the pre-training loss. Consequently, there is a set of minimizers $\bm{W}$ which differ only in those degenerate directions. Among these, the minimal-Frobenius-norm criterion forces all degenerate coordinates to be zero as any nonzero component in that subspace would increase $\|\bm{W}\|_{F}$. Consequently,
@@
For the second part of the proposition, the argument directly follows from the proof of Theorem 5.3. ∎

### v2

@@ Appendix B Proofs for Section 4
$={\bm{\beta}_{\text{TT}}}^{\top}\left[\bm{\Sigma}_{\bm{x}}\;-\;n\bm{\Sigma}_{\bm{x}}\,\bm{W}\,\bm{\Sigma}_{\bm{x}}\;-\;n\bm{\Sigma}_{\bm{x}}\bm{W}^{\top}\,\,\bm{\Sigma}_{\bm{x}}\;+\;n(n+1)\bm{\Sigma}_{\bm{x}}\bm{W}^{\top}\bm{\Sigma}_{\bm{x}}\bm{W}\bm{\Sigma}_{\bm{x}}+n\texttt{tr}(\bm{W}^{\top}\bm{\Sigma}_{\bm{x}}\bm{W}\bm{\Sigma}_{\bm{x}})\bm{\Sigma}_{\bm{x}}\right]{\bm{\beta}_{\text{TT}}}+\sigma^{2}n\texttt{tr}(\bm{W}^{\top}\bm{\Sigma}_{\bm{x}}\bm{W}\bm{\Sigma}_{\bm{x}})+\sigma^{2}.$
@@
Recall the standard result, which follows from Wick’s Theorem (Petersen and Pedersen, 2012)[Section 8.2.4] for a zero-mean Gaussian $\bm{x}\sim\mathcal{N}(\bm{0},\bm{\Sigma})$ and any symmetric matrix $\bm{A}$}:
@@
Hence, considering that we have $n$ diagonal terms and n(n1)n(n-1) off-diagonal terms, combining (i) + (ii) gives:
@@
${\cal{L}}(\bm{W})$ $={\bm{\beta}_{\text{TT}}}^{\top}\left[\bm{\Sigma}_{\bm{x}}\;-\;n\bm{\Sigma}_{\bm{x}}\,\bm{W}\,\bm{\Sigma}_{\bm{x}}\;-\;n\bm{\Sigma}_{\bm{x}}\bm{W}^{\top}\,\,\bm{\Sigma}_{\bm{x}}\;+\;n(n+1)\bm{\Sigma}_{\bm{x}}\bm{W}^{\top}\bm{\Sigma}_{\bm{x}}\bm{W}\bm{\Sigma}_{\bm{x}}+n\texttt{tr}(\bm{W}^{\top}\bm{\Sigma}_{\bm{x}}\bm{W}\bm{\Sigma}_{\bm{x}})\bm{\Sigma}_{\bm{x}}\right]{\bm{\beta}_{\text{TT}}}+\sigma^{2}n\texttt{tr}(\bm{W}^{\top}\bm{\Sigma}_{\bm{x}}\bm{W}\bm{\Sigma}_{\bm{x}})+\sigma^{2}.$
@@
We consider the noiseless ($\sigma^{2}=0$) and isotropic case where $\bm{\Sigma}_{\bm{x}}=\bm{\Sigma}_{\bm{\beta}}={\bm{I}}$. Throughout the proof, we omit lower order terms like +1+1 in $(d+1)$ or $2n$ in $(n^{2}+2n)$ as they’re negligible compared to the higher order of the same variable (see Remark˜4.3 for error discussion). This convention allows us to simplify the subsequent expressions.
@@
A key technical challenge while using the above update is that the expectation of $\Delta{\cal{L}}$ in (10) involves $8th$ moments of $\bm{X}_{\text{context}}$}. To overcome this problem, we will use the Gaussian approximation following Lemma B.1 by approximating ${\bm{X}_{\text{context}}}^{\top}{\bm{X}_{\text{context}}}\,{\bm{\beta}_{\text{TT}}}\approx n\bm{\Sigma}_{\bm{x}}\,{\bm{\beta}_{\text{TT}}}+\sqrt{n}\,\bm{\Sigma}_{\bm{x}}^{1/2}{\bm{g}}$ where ${\bm{g}}\sim\mathcal{N}(\bm{0},\|{\bm{\beta}_{\text{TT}}}\|^{2}\,{\bm{I}})$. This non-isotropic form can be obtained by factoring out $\bm{\Sigma}_{\bm{x}}^{1/2}$ and reducing the problem to the isotropic case from Lemma˜B.1. Also define $\Delta\bm{W}:={\bm{W}_{\text{TT}}}-\bm{W}^{*}$. Then, using the approximation and the fact that $\bm{\Sigma}_{\bm{x}}={\bm{I}}$, we can write:
@@
[…]
@@
$=-4\eta kn(n+1)\frac{nd}{(n+d+1)^{2}}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4}-4\eta^{2}\frac{k(k+d+1)n^{2}(n+1)}{(n+d+1)^{2}}\left[n\left(d^{2}-2d-3+nd\right)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{6}+\left((d+1)^{2}+n(d+2)\right)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{6}\right],$ (14)
@@
${\cal{L}}(\bm{W}^{*})-{\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$ $=4\eta k\frac{n^{2}d}{n+d+1}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4}-4\eta k\frac{n^{2}}{(n+d+1)^{2}}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4}-4\eta k\frac{n^{2}(n+1)d}{(n+d+1)^{2}}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4}$
$\quad-4\eta^{2}\frac{k(k+d+1)n^{2}}{(n+d+1)^{2}}\left[n(n+2)(d^{2}-2d-3+nd)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{6}+(n+d+1)\left((d+1)^{2}+n(d+2)\right)\|{\bm{\beta}_{\text{TT}}}\|_{2}^{6}\right].$
@@
Taking the derivative of this expression and setting to 0 yields that the optimal $\eta^{*}$ solution is approximately:
@@
See 4.4
@@
Using Theorem˜4.2, the new loss ${\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$ is approximately:
@@
$f^{\prime}(a)~=\;1~-\;3\,\frac{k}{k+d}a^{2}.$
Solving $f^{\prime}(a)=0$ yields $a_{\mathrm{crit}}=\sqrt{\;\frac{k+d}{\,3k\,}}$. This lies in $(0,1)$ when $\sqrt{\tfrac{k+d}{3k}}<1$, i.e. $k>\tfrac{d}{2}$. In that case, $f(a)$ has exactly one turning point in $(0,1)$, so $f$ increases on $(0,a_{\mathrm{crit}})$ and decreases on $(a_{\mathrm{crit}},1)$, making it non-monotonic. On the other hand, in the $k<\frac{d}{2}$ regime, one finds
$\sqrt{\;\frac{k+d}{\,3k\,}\;}~>\;1\quad\implies\quad a_{\mathrm{crit}}\notin(0,1),$
@@
$\operatorname{\mathbb{E}}_{{\bm{X}_{\text{train}}},{\bm{\xi}_{\text{train}}},{\bm{X}_{\text{context}}},{\bm{\xi}_{\text{context}}}}\left[{\bm{\beta}_{\text{TT}}}^{\top}\,n\,\bm{\Sigma}_{\bm{x}}\,\left({\bm{W}_{\text{TT}}}-\bm{W}^{*}\right)\,\bm{\Sigma}_{\bm{x}}\,{\bm{\beta}_{\text{TT}}}\right]$ $=2\eta n\,{\bm{\beta}_{\text{TT}}}^{\top}\operatorname{\mathbb{E}}_{{\bm{X}_{\text{train}}},{\bm{X}_{\text{context}}}}\left[{\bm{X}_{\text{train}}}^{\top}{\bm{X}_{\text{train}}}\,{\bm{\beta}_{\text{TT}}}\,{\bm{\beta}_{\text{TT}}}^{\top}{\bm{X}_{\text{context}}}^{\top}{\bm{X}_{\text{context}}}\right]{\bm{\beta}_{\text{TT}}}$
$=2\eta kn^{2}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{4},$ (20)
@@
[…]
@@
Setting the derivative to 0 and solving for $\eta$ gives:
@@
In the proportional $n,d$ regime and when $k/d\to\infty$, this gives an error of order $\mathcal{O}(n^{-1})$-matching that of the optimal $\bm{W}^{\text{opt}}$} from Lemma˜C.1:
@@
Considering that the initial loss of the weight matrix $\bm{W}=\bm{0}_{d\times d}$ is $\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}$ and the improvement given by Theorem˜4.5, the new loss after the test-time-training update is approximately:
@@
On the other hand, recall that by Proposition˜4.1, the loss of $\bm{W}^{*}$ before the update is $\frac{d+1}{n+d+1}\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}$. Combining this with the improvement given by Theorem˜4.2, the new loss after the test-time-training update is approximately:
@@ Appendix C Proofs for Section 5
$\nabla_{\bm{W}}{\cal{L}}(\bm{W})=-\,2n\,{\bm{\beta}_{\text{TT}}}\,{\bm{\beta}_{\text{TT}}}^{\top}+2n(n+1)\bm{W}\,{\bm{\beta}_{\text{TT}}}\,{\bm{\beta}_{\text{TT}}}^{\top}\;+\;2n\,{\bm{\beta}_{\text{TT}}}^{\top}{\bm{\beta}_{\text{TT}}}\bm{W}\;=\;0.$
@@
Hence, all other eigenvalues of $\bm{W}$ are 0. Using the eigendecomposition of $\bm{W}$, this uniquely specifies $\bm{W}$ as $c\cdot{\bm{\beta}_{\text{TT}}}\,{\bm{\beta}_{\text{TT}}}^{\top}$. Solving for $c$ yields the unique solution to $\nabla_{\bm{W}}{\cal{L}}(\bm{W})=0$:
@@
Consider the optimal pre-trained $\bm{W}^{*}$ matrix, which minimizes the population loss over all possible tasks sampled from $\bm{\Sigma}_{\bm{\beta}}$ in (1). Then, all eigenvalues of $\bar{\bm{W}}^{*}=\bm{\Sigma}_{\bm{x}}^{1/2}\bm{W}^{*}\bm{\Sigma}_{\bm{x}}^{1/2}$ lie in the interval $0,\frac{1}{n+1}]$.
@@
preserves the population risks in (1) and (3). More precisely, defining $\bar{{\cal{L}}}(\bar{\bm{W}})$ to be the population loss (3) evaluated under the transformed setting $(\bar{\bm{x}},\bar{{\bm{X}}},\bar{{\bm{\beta}}},\bar{\bm{W}})$, we have
@@
Let $n/d=\Theta(1)$ and $\sigma^{2}=0$. Suppose the covariance matrices are jointly diagonalizable by an orthogonal matrix ${\bm{Q}}\in\mathbb{R}^{d\times d}$, so that $\bm{\Sigma}_{\bm{x}}={\bm{Q}}{\bm{\Lambda}}_{\bm{x}}{\bm{Q}}^{\top}$ and $\bm{\Sigma}_{\bm{\beta}}={\bm{Q}}{\bm{\Lambda}}_{\beta}{\bm{Q}}^{\top}$. Let $\bm{W}$ be any matrix that’s also jointly diagonalizable by $\bm{Q}$} such that if $\bm{\Sigma}_{\bm{x}}^{1/2}\,\bm{W}\,\bm{\Sigma}_{\bm{x}}^{1/2}={\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}$, then all eigenvalues of $\bm{\Lambda}$} lie in the interval $0,\frac{1}{n+1}]$. Define ${\tilde{\bm{\beta}}_{\text{TT}}}:={\bm{Q}}^{\top}\bm{\Sigma}_{\bm{x}}^{1/2}{\bm{\beta}_{\text{TT}}}$, $A:={\tilde{\bm{\beta}}_{\text{TT}}}^{\top}\left({\bm{I}}-n{\bm{\Lambda}}\right)^{2}{\tilde{\bm{\beta}}_{\text{TT}}}$, and $B:=n\|{\tilde{\bm{\beta}}_{\text{TT}}}\|_{2}^{2}\texttt{tr}({\bm{\Lambda}}^{2})$. Then, the optimal step size that minimizes the population loss given in (3) after the test-time training update is
@@
As in previous theorems, we will use the Gaussian approximation following Lemma B.1 by approximating ${\bm{X}_{\text{context}}}^{\top}{\bm{X}_{\text{context}}}\,{\bm{\beta}_{\text{TT}}}\approx n\bm{\Sigma}_{\bm{x}}\,{\bm{\beta}_{\text{TT}}}+\sqrt{n}\,\bm{\Sigma}_{\bm{x}}^{1/2}{\bm{g}}$ where ${\bm{g}}\sim\mathcal{N}(\bm{0},\|{\bm{\beta}_{\text{TT}}}^{2}\|\,{\bm{I}})$. This non-isotropic form can be obtained by factoring out $\bm{\Sigma}_{\bm{x}}^{1/2}$ and reducing the problem to the isotropic case from Lemma˜B.1. Before going forward, let’s apply the covariance shift discussed in Lemma C.3, which transforms:
@@
$=2\eta{\bar{\bm{X}}_{\text{train}}}^{\top}{\bar{\bm{X}}_{\text{train}}}\left(n\,{\bar{\bm{\beta}}_{\text{TT}}}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}+\sqrt{n}\,{\bar{\bm{\beta}}_{\text{TT}}}{\bm{g}}^{\top}-n^{2}{\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}{\bar{\bm{\beta}}_{\text{TT}}}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}-n\sqrt{n}\,{\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}{\bar{\bm{\beta}}_{\text{TT}}}\,{\bm{g}}^{\top}-n\sqrt{n}\,{\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}{\bm{g}}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}-n{\bm{Q}}{\bm{\Lambda}}{\bm{Q}}^{\top}{\bm{g}}{\bm{g}}^{\top}\right)$
@@
Now, recall from Equation˜28 that the second-order difference is in the form:
@@
$\approx 4\eta kn^{2}\|{\bar{\bm{\beta}}_{\text{TT}}}\|_{2}^{2}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}\,{\bm{Q}}\left({\bm{I}}-n{\bm{\Lambda}}\right)^{2}{\bm{Q}}^{\top}{\bar{\bm{\beta}}_{\text{TT}}}-4\eta^{2}k(k+d+1)n^{2}\left[n^{3}\,\|{\bar{\bm{\beta}}_{\text{TT}}}\|_{2}^{6}\,\texttt{tr}({\bm{\Lambda}}^{2})+(n^{2}+d)\,\|{\bar{\bm{\beta}}_{\text{TT}}}\|_{2}^{4}\,{\bar{\bm{\beta}}_{\text{TT}}}^{\top}\,{\bm{Q}}\left({\bm{I}}-n{\bm{\Lambda}}\right)^{2}{\bm{Q}}^{\top}{\bar{\bm{\beta}}_{\text{TT}}}\right].$ (39)
This is a quadratic expression in $\eta$, and we can solve for the optimal $\eta^{*}$ by setting the derivative to 0. This way, the optimal step size is approximately:
@@
While Li et al. (2024) derives the solution form in the full-rank case, the same closed-form expression can be adapted when $\bm{\Sigma}_{\bm{\beta}}$ is low-rank. Indeed, if $\bm{\Sigma}_{\bm{\beta}}$ has zeros in its last drd-r diagonal entries, those coordinates do not appear in the pre-training loss. Consequently, there is a set of minimizers $\bm{W}$ which differ only in those degenerate directions. Among these, the minimal-Frobenius-norm criterion forces all degenerate coordinates to be zero as any nonzero component in that subspace would increase $\|\bm{W}\|_{F}$. Consequently,
@@
For the second part of the proposition, the argument directly follows from the proof of Theorem˜5.3. ∎